In the following series which comes next 0,5,22,57,?,205.Please provide explanation for this question.
Answer by stanbon(75887)

In the following series which comes next 0,5,22,57,?,205.Please provide explanation for this question.
----
My cubic regression program says
y = x^3 - 2x + 1
----
f(5) = 5^3-2*5+1 = 125-10+1 = 134
-------------------
Cheers,
Stan H.
===================